OpenStudy (anonymous):

Need help setting up two equations properly. However, I can solve the rest... just need help setting them up! :)

OpenStudy (anonymous):

1. Five less than twice the square a number, the result is 7 more than three times a number 2. A motorcycle can travel 705 miles on 19 gallons of gas. How many gallons of gas are needed for traveling 1253 miles?

OpenStudy (yamyam70):

1. ) 5 - (2x)^2 = 7 + 3x 2.) I don't know :D

OpenStudy (anonymous):

Lol, thanks for your help anyhow @yamyam70 :)

OpenStudy (yamyam70):

:)

OpenStudy (texaschic101):

1. 2(x^2) - 5 = 3x + 7 2. I know a proportion... 705/19 = 1253/x (705 miles to 19 gallons = 1253 miles to x gallons) you can solve this by cross multiplying and then solve for x.

OpenStudy (mathstudent55):

The best way to do 1. is to deal with each part of the sentence and build up the expression. "1. Five less than twice the square a number, the result is 7 more than three times a number" Let x = a number The "square of a number " is \(x^2\) "twice the square of a number is" \(2x^2\) "five less than twice the square of a number" is \(2x^2 - 5\) "three times a number" is 3x "7 more than three times a number" is 3x + 7 Now we equate both parts: \(2x^2 - 5 = 3x + 7\)

OpenStudy (mathstudent55):

@yamyam70 Be careful with "5 less than y." It is y - 5, not 5 - y.

OpenStudy (yamyam70):

oh yeah, thanks again @mathstudent55 :)

OpenStudy (texaschic101):

mathstudent....on the second one, do you solve it with a proportion ?

OpenStudy (mathstudent55):

@texaschic101 Yes, you did a great job, and your method of solving is good too.

OpenStudy (texaschic101):

thank you

OpenStudy (mathstudent55):

wlcm